your weekly paycheck is 30% less then your co worker your two paychecks total 740 find the amount of each paycheck
Answer by stanbon(75887)

your weekly paycheck is 30% less then your co worker.
your two paychecks total 740
find the amount of each paycheck
----
Let one of the paycheck be "x".
---
The other paycheck is "0.7x".
---
Equation:
x + 0.7x = 740
1.7x = 740
x = $435.29 (one of the paychecks)
0.7x = 0.7*435.29 = $304.71 (the other paycheck)
==================
Cheers,
Stan H.